OpenStudy (anonymous):

Where are the asymptotes of f(x) = tan 4x from x = 0 to x = pi over two?

OpenStudy (dtan5457):

Asymptotes of a tan function are found form the formula \[\theta=\frac{ n \pi }{ 2 }\]

OpenStudy (dtan5457):

theta=the value of b (in this case it's 4x)

OpenStudy (anonymous):

What does the n stand for?

OpenStudy (dtan5457):

n=any odd integer...so once you solve for x...n can represent many numbers. (all asymptotes)

OpenStudy (anonymous):

so it would be like \[4x = \frac{ 5\pi }{ 2 }\]

OpenStudy (dtan5457):

well 5pi could be one of the asymptotes. but basically you want \[x=\frac{ n \pi }{ 8 }\]

OpenStudy (anonymous):

Where did the 8 come from?

OpenStudy (dtan5457):

you multiplied out the 4x to the denominator.

OpenStudy (anonymous):

oh okay that makes more sense. But I am still confused of how to solve that equation to find the asymptotes

OpenStudy (dtan5457):

well, they give the domain from 0 to one period (pi/2) start putting odd numbers in that equation to see if it fits the domain.

OpenStudy (dtan5457):

from what i can see, there are two asymptotes

OpenStudy (dtan5457):

at least from 0 to pi/2

OpenStudy (dtan5457):

any trouble?

OpenStudy (anonymous):

Sorry my laptop is acting up, here are the answer choices: x = pi over 4, x = 3 pi over 4 x = pi over 8, x = 3 pi over 8 x = 0, x = pi over 4 x = pi over 2, x = 3 pi over 2

OpenStudy (dtan5457):

yup it's in there alright

OpenStudy (anonymous):

I am still confused on how to figure it out. I know to put in odd numbers, but should I put in the numbers from the answer choices?

OpenStudy (dtan5457):

Knowing that your domain is from 0 to pi/2 and that your asymptote finder equation is \[\frac{ n \pi }{ 8 }\] that basically means n can't be greater than 3 as pi/2=4pi/8

OpenStudy (dtan5457):

4 is a even number anyhow so it wouldn't work regardless

OpenStudy (anonymous):

So, the answer would be C then?

OpenStudy (dtan5457):

no your equation was npi/8 the denominator must be 8 i'm pretty sure i just gave it away lol

OpenStudy (anonymous):

Oh sorry I meant to say B. I'm not that good in math as you can tell

OpenStudy (dtan5457):

it's alright, you actually tried which is all that matters to me :)

OpenStudy (dtan5457):

anymore questions ?

OpenStudy (anonymous):

Yeah, there is one section that I always get wrong: Compare each of the functions shown below: f(x) graph of a downward facing parabola with vertex at 2, 3 g(x) = 2 cos(2x − π) + 2 h(x) x y −1 −7 0 −2 1 1 2 2 3 1 4 −2 5 −7 Which function has the largest maximum?

OpenStudy (dtan5457):

So...3 functions, and you want to know which the largest maximum value?

OpenStudy (anonymous):

Yeah, I tried figuring it out but I don't know if it is right

OpenStudy (dtan5457):

Do you know if it's asking for the max x value...or max y value..?

OpenStudy (anonymous):

It doesn't say if it is or not. But I assumed it meant maximum y value

OpenStudy (dtan5457):

Alright, then I think the max y value should be fairly clear in functions f(x) and h(x), can you tell me what they are ?

OpenStudy (anonymous):

The maximum for f(x) is 3 and the maximum for h(x) would be 2

OpenStudy (dtan5457):

Now the slightly harder one.. do you know the max/min of g(x) = 2 cos(2x − π) + 2?

OpenStudy (anonymous):

I have no idea

OpenStudy (dtan5457):

In the function g(x) = 2 cos(2x − π) + 2 there is a amplitude of 2, but also a vertical shift of +2 your max would be \[k+\left| a \right|\] min would be \[k-\left| a \right|\]

OpenStudy (dtan5457):

k represents the vertical shift (+2) a=amplitude

OpenStudy (dtan5457):

this is for the y value of course

OpenStudy (anonymous):

So the maximum y value would be 4?

OpenStudy (dtan5457):

yep

OpenStudy (anonymous):

Okay I get it now thank you!

OpenStudy (dtan5457):

yw. feel free to fan and medal, and tag me onto more questions :)